Why the velocity of light  in vacuum is independent of  wave length ?

Grade:12

3 Answers

India VK
10 years ago

Speed of any wave does not depend on the wavelength. It depends only on the elastic and inertia properties of the medium.

srinu kummari
10 years ago

In vaccum if frequency increases wavelenth decreases, if wavelenth increases frequency decreases so velacity in vaccum is independent of wavelenth

Naveen
10 years ago

na= refractive index of air   = 1

nb=refractive index of vacume = 0

va=velocity of light in air  =3*10^8m/s

vb=velocity of light in vacume

 

we know that na/nb= vb/va

then vb = 3*10^8m/s

 

fa  =  va/lmdab  ;  fb  =  vb/lmdab

by this fa=fb&va=vb

then lmda a=lmada b

so that the velocity of light  in vacuum is independent of  wave length